# Social Gesture Recognition in spHRI: Leveraging Fabric-Based Tactile Sensing on Humanoid Robots

### v2

@@ I INTRODUCTION
In this paper, we introduce a system for social gesture recognition that incorporates a fabric-based, large-scale tactile sensor attached to the arms of a humanoid robot. To classify gestures, we built a social gesture dataset by collecting data from various participants and extracting temporal features for analysis. Our sensor was machine-knitted and designed to maintain accuracy even when the robot is in motion. The flexibility of the machine-knitted sensor enables easy customization to fit different shapes, while real-world data collection on an actual humanoid robot offers valuable insights into how humans interact with robots through touch. We believe that developing better systems for gesture recognition will allow for more natural and enjoyable social-physical human-robot interaction.
@@ II-B Social Physical Human Robot Interaction
Humanoid robots can assist in various settings, including healthcare [31], education [32], and companionship [31]. So we implemented our system on Pollen Robotics Reachy 2023 (Fig. 1). The robot’s arm has seven degrees of freedom and they tend to mirror the arms of a human. The robot also has more of a cartoon looking face making the robot have more friendly look.
@@ II-D Social Gesture Recognition
Enabling gesture recognition in robots could provide a way for them to understand messages communicated through touch. Humans are capable of conveying and interpreting various emotions through touch [42], and researchers have been exploring how humans use touch to express emotions to robots. Yohanan and MacLean, for example, examined the types of gestures people use to convey specific emotions [43]. They also created a gesture touch dictionary based on human-animal and human-human touch, as well as affective touch [43]. Understanding touch gestures is a promising avenue for creating an additional communication modality between humans and robots, which has led to significant research in gesture recognition. One notable resource in this area is the CoST dataset, which provides valuable pressure data on various social gestures performed on a sensor [44]. Some research has focused on feature extraction from the CoST dataset, including spatiotemporal fusion features, for classification tasks [45]. While the CoST dataset offers a solid foundation by collecting diverse pressure data from different individuals performing various social gestures, it does not involve humanoid robots, which may elicit different gestures. Additionally, the sensor used in the CoST dataset differs from the one we are employing, meaning that we cannot directly use their dataset. Other studies have also contributed to the field of social gesture recognition [9].
@@ III Sensor Design and Fabrication
To increase sensing coverage, we attached the sensor to the robot in a way which prevents the fabric from moving while the arm is in motion. Our sensor is composed of two components. One is affixed to the upper arm of the robot, while the other is attached to the lower arm (Fig. 2). In the figure, the area where the black strips span is the sensing area. The blue portion which is covering the joints is there for aesthetic purposes only. This means that the elbow region of the arm is not covered by sensors. This area of coverage is to be done in future works. EZlifego double sided was used to fix the sensor on to the robot. Each sensing part of the sensor (the blue and purple boxes in Fig. 2 (B)) was attached to only one link, to prevent the sensor from giving readings when the robot’s arm was in motion Fig. 3. The robot has holes in the arm (Fig. 2 (A)) which needed to be addressed. Using duct tape, the holes where covered, but this caused a decrease in sensor performance (view the experiments section).
@@ IV-B Data Prepossessing and Feature Extraction
inline,color=red!40]The content is not clear. Please be more specific of the method you used, and explain the details. We preprocessed and extracted temporal features from our 3D spatial temporal data. We removed any frames recorded before the initial contact on the sensor. This was necessary because sensor output was recorded before interaction occurred. The feature extracted was the number of taxels activated per frame. An activated taxel is defined by having a voltage reading which is larger than zero. For each frame, we summed the number of taxels that were activated. This resulted in an array of numbers representing the number of taxels activated for every frame. Each gesture varied in frame length so the data was either padded or clipped to be 150 frames. The number of taxels tended to be unique for each gesture. Other features, like pressure information or more spatial features did not yield a higher accuracy.
@@ V-B2 Data Collection
To create our dataset we collected gesture data from various participants. Participants watched a video of the gesture being performed on a teddy bear. A stuffed animal was used for the visualization instead of using the sensor attached to the robot arm to avoid repeating exact hand positions seen in the video. Videos were used instead of displaying text to the participant to ensure each person interpreted the gesture properly. We noticed during some preliminary trials that the participants would interpret and perform the gestures differently.
@@ V-B3 Gesture Recognition Results
Our method was able to achieve an accuracy of 81.16% which is higher than our highest baseline, 76.78%, but still confused gestures for one another. Gestures "hit" and "poke" were often confused with each other, and "shake" and "grab" frequently led to similar misclassification issues as well. This is because there are times in which the signal for hit and poke are both the same. This happens when a participant would hit the sensor but they would have a bone or knuckle which is outstretched more than the rest of their fist. This means that only one or two taxels will get activated for around the same amount of time and with a similar force as poke.

### v3

@@ I INTRODUCTION
In this paper, we introduce a system for social gesture recognition that incorporates a fabric-based, large-scale tactile sensor attached to the arms of a humanoid robot. To classify gestures, we built a social gesture dataset by collecting data from various participants and extracting temporal features for analysis. Our sensor was machine-knitted and designed to maintain accuracy even when the robot is in motion. The flexibility of the machine-knitted sensor enables easy customization to fit different shapes, while data collection on an actual humanoid robot offers insights into how humans interact with robots through touch. We believe that developing better systems for gesture recognition will allow for more natural and enjoyable social-physical human-robot interaction.
@@ II-B Social Physical Human Robot Interaction
Humanoid robots can assist in various settings, including healthcare [31], education [32], and companionship [31]. So we implemented our system on Pollen Robotics Reachy 2023 (Fig. 1). The robot’s arm has seven degrees of freedom and they tend to mirror the arms of a human. The robot also has more of a cartoon looking face making the robot have a more friendly look.
@@ II-D Social Gesture Recognition
Enabling gesture recognition in robots offers a promising way for them to interpret messages communicated through touch. Humans naturally convey and interpret various emotions through tactile interactions [42], and researchers have been investigating how these emotional expressions translate to human-robot touch communication [43]. For instance, Yohanan and MacLean examined the types of gestures people use to express specific emotions [44]. They also developed a gesture touch dictionary inspired by human-animal and human-human touch, as well as affective touch patterns [44]. Understanding touch gestures presents an opportunity to create an additional communication modality between humans and robots, driving research into gesture recognition. One notable resource is the CoST dataset, which provides pressure data from various social gestures performed on a sensor [45]. Some studies have focused on extracting features from such gesture datasets, including spatiotemporal characteristics, for classification tasks [46, 9]. Recent work has also explored incorporating shear forces into touch gesture recognition, resulting in improved classification accuracy [47].
Although the CoST dataset offers a valuable foundation by capturing diverse pressure data from individuals performing different social gestures, it does not involve humanoid robots, which may evoke different gesture behaviors. Furthermore, the sensor used in the CoST dataset differs from the one employed in our research, preventing direct application of their dataset.
@@ III Sensor Design and Fabrication
We attached the sensor to the robot in a way which prevents the fabric from moving while the arm is in motion. Our sensor is composed of two components. One is affixed to the upper arm of the robot, while the other is attached to the lower arm (Fig. 2). In the figure, the area where the black strips span is the sensing area. The blue portion which is covering the joints is there for aesthetic purposes only. This means that the elbow region of the arm is not covered by sensors. This area of coverage is to be done in future works. EZlifego double sided was used to fix the sensor on to the robot. Each sensing part of the sensor (the blue and purple boxes in Fig. 2 (B)) was attached to only one link, to prevent the sensor from giving readings when the robot’s arm was in motion Fig. 3. The robot has holes in the arm (Fig. 2 (A)) which needed to be addressed. Using duct tape, the holes where covered, but this caused a decrease in sensor performance.
@@ IV-B Data Prepossessing and Feature Extraction
inline,color=red!40]The content is not clear. Please be more specific of the method you used, and explain the details. We preprocessed and extracted temporal features from our 3D spatial temporal data. We removed any frames recorded before the initial contact on the sensor. This was necessary because sensor output was recorded before interaction occurred. The feature extracted was the number of taxels activated per frame. An activated taxel is defined by having a digital reading which is larger than 10. For each frame, we summed the number of taxels that were activated. This resulted in an array of numbers representing the number of taxels activated for every frame. Each gesture varied in frame length so the data was either padded or clipped to be 150 frames. The number of taxels tended to be unique for each gesture. Other features, like pressure information or more spatial features did not yield a higher accuracy.
@@ V-B2 Data Collection
To create our dataset we collected gesture data from various participants. Participants watched a video of the gesture being performed on a teddy bear. A stuffed animal was used for the visualization instead of using the sensor attached to the robot arm to avoid repeating exact hand positions seen in the video. Videos were used instead of displaying text to the participant to ensure each person interpreted the gesture properly. We noticed during some preliminary trials that the participants would interpret and perform the gestures differently which is why we included the demonstration.
@@ V-B3 Gesture Recognition Results
Our method was able to achieve an accuracy of 81.16% which is higher than our highest baseline, 76.78%, but still confused gestures for one another. Gestures "hit" and "poke" were often confused with each other, and "shake" and "grab" frequently led to similar misclassification issues as well. This is because there are times in which the signal for hit and poke are both the same. This happens when a participant would hit the sensor but they would have a bone or knuckle which is outstretched more than the rest of their fist. This means that only one or two taxels will get activated for around the same amount of time and with a similar force as poke. These results seem promising, however more data should be collected on how the system performs when people perform the same gesture differently.